Alvin paddled for 5 hours with a 3​-km/h current to reach a campsite. The return trip against the same current took 8 hours.
What is the speed of the boat in still​ water?

Answer by Theo(13342)

rate * time = distance.

let r = rate and t = time and d = distance

going downstream, the formula becomes (r+3) * 5 = d
coming back upstream, the formula becomes (r-3) * 8 = d

the distance downstream is the same as the distance upstream, so (r+3) * 5 = (r-3) * 8

simplify to get 5r + 15 = 8r - 24

subtract 5r from both sides of this equation and add 24 to both sides of this equation to get 39 = 3r

solve for r to get r = 39/3 = 13

he paddles at 13 kilometers per hour.

going downstream, the stream adds 3 kilometers per hours to this to get 16 kilometers per hour for 5 hours = 80 kilometers.

coming back upstream, the stream subtracts 3 kilometers from this to get 10 kilometers per hour for 8 hours = 80 kilometers.

the distance downstream and upstream is the same at 80 kilometers.

the speed of the boat is 13 kilometers per hour